# The Nonlinear Filter Model of Stream Cipher Redivivus

### v2

@@ Abstract
The nonlinear filter model is an old and well understood approach to the design of secure stream ciphers. Extensive research over several decades has shown how to attack stream ciphers based on this model and has identified the security properties required of the Boolean function used as the filtering function to resist such attacks. This led to the problem of constructing Boolean functions which provide adequate security and at the same time are efficient to implement. Unfortunately, over the last two decades no good solutions to this problem appeared in the literature. The lack of good solutions has effectively led to nonlinear filter model becoming more or less obsolete. This is a big loss to the cryptographic design toolkit, since the great advantages of the nonlinear filter model are its simplicity, well understood security and the potential to provide low cost solutions for hardware oriented stream ciphers. In this paper we construct balanced functions on an odd number $n\geq 5$ of variables with the following provable properties: linear bias equal to $2^{-\lfloor n/2\rfloor-1}$, algebraic degree equal to $2^{\lfloor\log_{2}\lfloor n/2\rfloor\rfloor}$, algebraic immunity at least $\lceil(n-1)/4\rceil$, fast algebraic immunity at least $1+\lceil(n-1)/4\rceil$, and can be implemented using $O(n)$ NAND gates. The functions are obtained from a simple modification of the well known class of Maiorana-McFarland bent functions. By appropriately choosing $n$ and the length $L$ of the linear feedback shift register, we show that it is possible to obtain examples of stream ciphers which are provably $\kappa$-bit secure against well known classes of attacks for various values of $\kappa$. We provide concrete proposals for $\kappa=80,128,160,192,224$ and $256$. For the $80$-bit, $128$-bit, and the $256$-bit security levels, the circuits for the corresponding stream ciphers require about 1743.5, 2771.5, and 5607.5 NAND gates respectively. For the $80$-bit and the $128$-bit security levels, the gate count estimates compare quite well to the famous ciphers Trivium and Grain-128a respectively, while for the $256$-bit security level, we do not know of any other stream cipher design which has such a low gate count.
@@ 1 Introduction
The nonlinear filter model of stream ciphers is several decades old; one may note that the model was extensively discussed in the book by Rueppel [37] which was published in the mid-1980s. The nonlinear filter model consists of two components, namely a linear feedback shift register (LFSR) and a Boolean function $f$ which is applied to a subset of the bits of the LFSR. At each clock, $f$ is applied to the present state of the LFSR to produce a single keystream bit and simultaneously the LFSR also moves to the next state. LFSRs are very efficient to implement, particularly in hardware. So the implementation efficiency of the nonlinear filter model is essentially determined by the efficiency of implementation of $f$.
@@
Extensive research carried out over the last few decades has shown several approaches to cryptanalysing the nonlinear filter model of stream ciphers. The initial line of attack was based upon determining the linear complexity of the produced keystream. For an LFSR of length $L$ and a Boolean function of algebraic degree $d$, under certain reasonable and easy-to-ensure conditions, the linear complexity of the keystream is known to be at least (${L\choose d}$ (see [37]). Using large enough values of $L$ and $d$, linear complexity based attacks can be made infeasible. The second phase of attacks consisted of various kinds of (fast) correlation attacks. Starting with the first such attack in 1985 [41], a long line of papers [35, 23, 17, 28, 29, 15, 12, 30, 11, 16, 42, 45, 44, 31, 32] explored various avenues for mounting correlation attacks. Surveys of some of the older attacks appear in [9, 10, 33, 2]. The resistance to correlation attacks is mainly determined by the linear bias $\varepsilon$ of the Boolean function $f$. The linear bias is determined by the nonlinearity of the function $f$; the higher the nonlinearity, the lower the linear bias. The third phase of attacks started in 2003 with the publication of the algebraic attack [19] and was soon followed by the publication of the fast algebraic attack [18]. Resistance to these attacks requires the function $f$ to possess high (fast) algebraic immunity.
@@
The various attacks mentioned above have posed the following design challenge for a Boolean function to be used in the nonlinear filter model of stream ciphers. Construct balanced Boolean functions which achieve a good combination of nonlinearity and algebraic resistance and are also very efficient to implement. Unfortunately, since the time algebraic attacks were proposed in the early-2000s, no good solutions to the design challenge for Boolean functions have appeared in the literature. A consequence of not being able to find good solutions to the design challenge is that the nonlinear filter model of stream ciphers became obsolete. This is somewhat unfortunate since the model is very old, well studied with well understood security, and the potential to provide low gate count solutions in hardware.
@@
In this paper, we revisit the above mentioned design problem for Boolean functions towards the goal of reviving the nonlinear filter model. Bent functions [36] are a very well studied class of Boolean functions. They exist for even number of variables and provide the highest possible nonlinearity. However, their use in cryptography is not clear. The introduction of Chapter 6 of [13] summarises the state-of-the-art as follows: “we do not know an efficient construction using bent functions which would provide Boolean functions having all the necessary features for being used in stream ciphers.” Even though this is discouraging, nonetheless, we start our construction from bent functions.
@@
The well known Maiorana-McFarland class of bent functions is defined as follows. For $m\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $m$ variables. Then a $2m$-variable Maiorana-McFarland bent function is defined to be $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})$, where $\pi$ is a bijection from $m$-bit strings to $m$-bit strings and $h$ is any $m$-variable Boolean function. It is well-known that the nonlinearity of Maiorana-McFarland functions does not depend on the choice of $h$ nor on that of the permutation $\pi$. Bad choices of $\pi$ and $h$ (for example, $\pi$ to be the identity permutation and $h$ to be a constant function) provide functions whose algebraic immunity is low. On the bright side, we observe that it may be possible to improve the algebraic resistance by properly choosing $\pi$ and $h$. Assume that $\pi$ is chosen to be the identity permutation. It is known [20] that the majority function possesses the best possible algebraic immunity. Motivated by this fact, we choose $h$ to be the majority function on $m$ variables. We prove that the resulting bent function on $2m$ variables has algebraic immunity at least $\lceil m/2\rceil$, and hence fast algebraic immunity at least $1+\lceil m/2\rceil$. On the implementation aspect, we show that the majority function can be computed using $O(m)$ NAND gates. So the obtained bent function has maximum nonlinearity, good algebraic immunity and at the same time is quite efficient to implement. The algebraic immunity is not the maximum possible, but due to the implementation efficiency, it is possible to increase the number of variables to achieve the desired level of algebraic resistance. One small problem is that a bent function is not balanced. This problem is easily rectified by XORing a new variable to the bent function, obtaining a function on an odd number of variables. This modification does not affect the linear bias or the algebraic resistance and requires only one extra XOR gate for implementation. A novelty of our work is the observation and the proof that choosing $h$ to be the majority function improves the algebraic immunity. While Maiorana-McFarland bent functions have been extensively studied in the literature, this simple observation has escaped the notice of earlier researchers.
@@
We perform a detailed concrete security analysis of some of the well known attacks on the nonlinear filter model. As the outcome of this analysis, for various security levels, we provide concrete proposals for stream ciphers based on the nonlinear filter model using the Boolean functions described above as the filtering functions. A strong point in favour of these proposals is that at the appropriate security levels they provide provable assurance against well known classes of attacks. Further, we provide concrete gate count estimates for the entire circuit to implement the stream ciphers. For the 80-bit, $128$-bit, 160-bit, 192-bit, 224-bit and the $256$-bit security levels, we obtain gate count estimates of 1743.5, 2771.5, 3520.5, 4188.5, 4854.5, and 5607.5 NAND gates respectively. The gate count estimates for the 80-bit and the 128-bit security levels compare quite well [^1] Following [1] we estimated 8 NAND gates for a flip-flop, whereas Trivium estimated 12 NAND gates for a flip-flop. Using 12 NAND gates for a flip-flop, our proposal at the 80-bit security level requires about 2395.5 NAND gates, whereas Trivium requires about 3488 NAND gates. with famous ciphers such as Trivium [8] and Grain-128a [1] which offer 80-bit and 128-bit security respectively. For the other security levels, we are not aware of other stream ciphers which have such low gate counts. So our revival of the nonlinear filter model of stream ciphers leads to concrete proposals which offer a combination of both provable security against well known classes of attacks at a desired level of security and also low gate count.
@@
We note that there are some old (prior to the advent of algebraic attacks) works [39, 40, 24] on efficient implementation of Boolean functions on a large number of variables targeted towards the nonlinear combiner model of stream ciphers. In the context of the nonlinear filter model, however, we are not aware of any previous work on the design and implementation of Boolean functions on a large number of variables. So this aspect is also new to the present work.
@@ 4 Concrete Proposals
We revisit the filter generator model where the state of an LFSR of length $L$ is filtered using a Boolean function. By $\mathcal{S}(L,m)$, we denote the class of stream ciphers obtained from an LFSR of length $L$ with a primitive connection polynomial, and with $f_{2m+1}$ as the filtering function. For the $\kappa$-bit security level, we assume that the stream cipher supports a $\kappa$-bit secret key and a $\kappa$-bit initialisation vector (IV), and so $L\geq 2\kappa$. Suppose that the state $(s_{L-1},\ldots,s_{0})$ of the LFSR leads to the next state $(s_{L},\ldots,s_{1})$, i.e. the state is right shifted once and the next bit $s_{L}$ becomes the leftmost bit. The initial state of the LFSR is constructed as $(k_{\kappa-1},\ldots,k_{0},v_{\kappa-1},\ldots,v_{0},b_{1},\ldots,b_{L-2\kappa})$, where $(k_{\kappa-1},\ldots,k_{0})$ is the key, $(v_{\kappa-1},\ldots,v_{0})$ is the IV, and $(b_{1},\ldots,b_{L-2\kappa})\neq\mathbf{0}_{L-2\kappa}$ is an arbitrary constant bit string. We next specify which state bits of the LFSR are fed into the filtering function. From the state $(s_{L-1},\ldots,s_{0})$ of the LFSR, the value of $W$ in $f_{2m+1}(W,\mathbf{X},\mathbf{Y})$ is the state bit $s_{L-\kappa+m}$, the values of the variables in $\mathbf{X}$ are the state bits $s_{L-\kappa+m-1},\ldots,s_{L-\kappa}$, and the values of the variables in $\mathbf{Y}$ are the state bits $s_{L-2\kappa+m-1},\ldots,s_{L-2\kappa}$. We assume that the stream cipher goes through an initialisation phase of $2\kappa$ steps, where no keystream bit is produced, and instead the keystream bit is fed back by XORing it with the next bit of the LFSR. So after the initialisation phase the leftmost $2\kappa$ bits of the LFSR are complex nonlinear functions of the key and the IV. Keystream generation starts after the initialisation phase.
@@
The parameters $L$ and $m$ are the design parameters of the stream cipher. We consider two basic attack parameters, namely the number $N$ of keystream bits that is required for a successful attack, and the time complexity $T$ of the attack. To ensure security at level $\kappa$ any attack should require $T>2^{\kappa}$. The condition $L\geq 2\kappa$ ensures that the size of the state is at least twice the size of the key. This prevents certain (theoretical) time/memory trade-off attacks [27]. We assume that at most $2^{B}$ keystream bits are generated from a single key and IV pair. So $N$ is at most $2^{B}$, as otherwise the attack cannot be mounted. Further, a basic condition is that $N\leq T$, i.e. some operation is performed on each keystream bit that is required. So for $\kappa$-bit security using $2^{B}$ keystream bits, we have $B\leq\kappa$. In the analysis below, our goal is to determine values of $L$ and $m$ such that the stream cipher is not vulnerable to the known types of attacks. For this concrete analysis, we ignore small constants appearing in the expressions for $N$ and $T$ (i.e. we consider these constants to be 1) and focus only on the exponential components of these quantities.

### v3

@@ Abstract
The nonlinear filter model is an old and well understood approach to the design of secure stream ciphers. Extensive research over several decades has shown how to attack stream ciphers based on this model and has identified the security properties required of the Boolean function used as the filtering function to resist such attacks. This led to the problem of constructing Boolean functions which provide adequate security and at the same time are efficient to implement. Unfortunately, over the last two decades no good solutions to this problem appeared in the literature. The lack of good solutions has effectively led to nonlinear filter model becoming more or less obsolete. This is a big loss to the cryptographic design toolkit, since the great advantage of the nonlinear filter model lies, beyond its simplicity and its ability to provide low-cost solutions for hardware-oriented stream ciphers, in the accumulated knowledge about the security requirements for tap positions and the filter function, which gives great confidence in its security when all criteria are met. In this paper we construct balanced functions on an odd number $n\geq 5$ of variables with the following provable properties: linear bias equal to $2^{-\lfloor n/2\rfloor-1}$, algebraic degree equal to $2^{\lfloor\log_{2}\lfloor n/2\rfloor\rfloor}$, algebraic immunity at least $\lceil(n-1)/4\rceil$, fast algebraic immunity at least $1+\lceil(n-1)/4\rceil$, and the functions can be implemented using $O(n)$ NAND gates. The functions are obtained from a simple modification of the well known class of Maiorana-McFarland bent functions. Due to the efficiency of implementation, for any target security level, we can construct efficient implementable functions which provide the required level of resistance to fast algebraic and fast correlation attacks. Previously known efficiently implementable functions have too large linear bias making them inappropriate even with a large number of variables. By appropriately choosing $n$ and the length $L$ of the linear feedback shift register, we show that it is possible to obtain examples of stream ciphers which are provably $\kappa$-bit secure against well known classes of attacks for various values of $\kappa$. We provide concrete proposals for $\kappa=80,128,160,192,224$ and $256$ using LFSRs of lengths 163, 257, 331, 389, 449, 521 and filtering functions on 75, 119, 143, 175, 203 and 231 variables. For the $80$-bit, $128$-bit, and the $256$-bit security levels, the circuits for the corresponding stream ciphers require about 1743.5, 2771.5, and 5607.5 NAND gates respectively. For the $80$-bit and the $128$-bit security levels, the gate count estimates compare quite well to the famous ciphers Trivium and Grain-128a respectively, while for the $256$-bit security level, we do not know of any other stream cipher design which has such a low gate count.
@@ 1 Introduction
The nonlinear filter model of stream ciphers is several decades old; one may note that the model was extensively discussed in the book by Rueppel [48] which was published in the mid-1980s. The nonlinear filter model consists of two components, namely a linear feedback shift register (LFSR) and a Boolean function $f$ which is applied to a subset of the bits of the LFSR at fixed positions (called tap positions). At each clock, $f$ is applied to the present state of the LFSR to produce a single keystream bit and simultaneously the LFSR also moves to the next state. LFSRs are very efficient to implement, particularly in hardware. So the implementation efficiency of the nonlinear filter model is essentially determined by the efficiency of implementation of $f$.
@@
Extensive research carried out over the last few decades has shown several approaches to cryptanalysing the nonlinear filter model of stream ciphers. The initial line of attack was based upon determining the linear complexity of the produced keystream. For an LFSR of length $L$ and a Boolean function of algebraic degree $d$, under certain reasonable and easy-to-ensure conditions, the linear complexity of the keystream is known to be at least (${L\choose d}$ (see [48]). Using large enough values of $L$ and $d$, linear complexity based attacks can be made infeasible. The second phase of attacks consisted of various kinds of (fast) correlation attacks. Starting with the first such attack in 1985 [53] (which was efficient on another model, the nonlinear combiner model), a long line of papers [52, 44, 29, 3, 21, 35, 36, 19, 13, 37, 12, 38, 20, 54, 60, 59, 39, 40] explored various avenues for mounting correlation attacks. Surveys of some of the older attacks appear in [10, 11, 42, 2]. Fast correlation attacks are based on affine approximation and apply to the nonlinear filter (as well as the nonlinear combiner) model. The resistance to fast correlation attacks is mainly determined by the linear bias $\varepsilon$ of the Boolean function $f$. The linear bias is determined by the nonlinearity of the function $f$; the higher the nonlinearity, the lower the linear bias. The third phase of attacks started in 2003 with the publication of the algebraic attack [23] and was soon followed by the publication of the fast algebraic attack [22]. Resistance to these attacks requires the function $f$ to possess high (fast) algebraic immunity.
@@
The various attacks mentioned above have posed the following design challenge for a Boolean function to be used in the nonlinear filter model of stream ciphers. Construct balanced Boolean functions which achieve a good combination of nonlinearity and algebraic resistance and are also very efficient to implement. Unfortunately, since the time algebraic attacks were proposed in the early-2000s, no good solutions to the design challenge for Boolean functions have appeared in the literature (some Boolean functions satisfy all the necessary cryptographic criteria, but are too heavy and slow to compute [17], and some others are fast to compute but possess insufficient nonlinearity [55, 57, 15, 41]). A consequence of not being able to find good solutions to the design challenge is that the nonlinear filter model of stream ciphers became obsolete. This is somewhat unfortunate since the model is very old, well studied with well understood security, and the potential to provide low gate count solutions in hardware.
A class of guess-then-determine attacks are known against nonlinear filter model. The inversion attack [27] is the first such attack. Subsequently, a number of papers have developed the idea into the generalised inversion attack, the filter state guessing attack, and the generalised filter state guessing attack [27, 28, 33, 46, 58]. Recommendations for resistance to the state guessing attack do not specify conditions on the filtering function. Rather, the recommendations specify conditions on the tap positions, i.e. the positions of the LFSR that are fed as input to the filtering function. We note that such recommendations implicity assume that the number of variables of the filtering function is much smaller than the desired security level.
@@
In this paper, we revisit the above mentioned design problem for Boolean functions towards the goal of reviving the nonlinear filter model. Bent functions [47] are a very well studied class of Boolean functions. They exist for even number of variables and provide the highest possible nonlinearity. However, their use in cryptography cannot be direct, since they are unbalanced, and even if we modify them into balanced functions (as Dobbertin [25] proposed) that use is not clear, after the invention of fast algebraic attacks. The introduction of Chapter 6 of [14] summarises the state-of-the-art as follows: “we do not know an efficient construction using bent functions which would provide Boolean functions having all the necessary features for being used in stream ciphers.” Note that a result from [56] seemed to imply that it was impossible to obtain a Boolean function having a good resistance to fast algebraic attacks by modifying a bent function into a balanced function. This result happened to be incorrect as shown in [14] (see Theorem 22 which corrects the result, and the few lines before it). Even after this correction, the problem of building good cryptographic functions from bent functions seemed hard. In the present paper we provide a concrete solution to this problem. We even show that it is possible to start with a bent function obtained from the very basic Maiorana-McFarland construction.
@@
The well known Maiorana-McFarland class of bent functions is defined as follows. For $m\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $m$ variables. Then a $2m$-variable Maiorana-McFarland bent function is defined to map $(\mathbf{X},\mathbf{Y})$ to $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})$, where $\pi$ is a bijection from $m$-bit strings to $m$-bit strings and $h$ is any $m$-variable Boolean function. It is well-known that the nonlinearity of Maiorana-McFarland functions does not depend on the choice of $h$ nor on that of the permutation $\pi$. Bad choices of both $\pi$ and $h$ (for example, $\pi$ to be the identity permutation and $h$ to be a constant function) provide functions whose algebraic immunity is low. On the bright side, we observe that it may be possible to improve the algebraic resistance by properly choosing $h$. Assume that $\pi$ is chosen to be the identity permutation. It is known [24] that the majority function possesses the best possible algebraic immunity. Motivated by this fact, we choose $h$ to be the majority function on $m$ variables. We prove that the resulting bent function on $2m$ variables has algebraic immunity at least $\lceil m/2\rceil$, and hence fast algebraic immunity at least $1+\lceil m/2\rceil$. On the implementation aspect, we show that the majority function can be computed using $O(m)$ NAND gates. So the obtained bent function has maximum nonlinearity, sufficient (fast) algebraic immunity when the number of variables is large enough (to protect against fast algebraic attacks at a specified security level), and at the same time is quite efficient to implement. Indeed, the (fast) algebraic immunity is not the maximum possible, but due to the implementation efficiency, it is possible to increase the number of variables to achieve the desired level of algebraic resistance. A novelty of our work is the observation and the proof that choosing $h$ to be the majority function improves the algebraic immunity. While Maiorana-McFarland bent functions have been extensively studied in the literature, this simple observation has escaped the notice of earlier researchers.
One problem is that a bent function is not balanced. This problem is easily rectified by XORing a new variable to the bent function, obtaining a function on an odd number of variables. This modification requires only one extra XOR gate for implementation. In terms of security, the modification does not change the linear bias. We prove that the algebraic immunity of the modified function is at least the algebraic immunity of the bent function, and hence the fast algebraic immunity is at least one plus the algebraic immunity of the bent function. So the algebraic resistance of the modified function is essentially the same as that of the bent function. A consequence of unbalancedness of the filtering function is that the keystream sequence is unbalanced, which leads to a distinguishing attack. The simple modification of XORing an extra variable prevents such a distinguishing attack. The resistance to fast correlation attacks and fast algebraic attacks are ensured by the bent function (on an appropriate number of variables). A positive aspect of XORing a new variable is that it prevents (see Theorem 2 of [27]) certain kinds of information leakage which was identified in [3]. On the negative side, filter functions of the form $W+g(\mathbf{Z})$ were shown to be susceptible to inversion attack [27]. We prevent this attack by choosing the gap between the tap positions to be sufficiently large; this countermeasure was already proposed in [27]. More generally, we show that due to the fact that we use filter functions on a large number of variables, the various kinds of state guessing attacks [28, 33, 46, 58] do not apply to our proposals.
@@
We perform a detailed concrete security analysis of some of the well known attacks on the nonlinear filter model. As the outcome of this analysis, for various security levels, we provide concrete proposals for stream ciphers based on the nonlinear filter model using the Boolean functions described above as the filtering functions. A strong point in favour of these proposals is that at the appropriate security levels they provide provable assurance against well known classes of attacks. Further, we provide concrete gate count estimates for the entire circuit to implement the stream ciphers. For the 80-bit, 128-bit, 160-bit, 192-bit, 224-bit and the $256$-bit security levels, we propose using LFSRs of lengths 163, 257, 331, 389, 449, 521 and filtering functions on 75, 119, 143, 175, 203 and 231 variables respectively. The gate count estimates for the 80-bit, 128-bit, 160-bit, 192-bit, 224-bit and the $256$-bit security levels are 1743.5, 2771.5, 3520.5, 4188.5, 4854.5, and 5607.5 NAND gates respectively. The gate count estimates for the 80-bit and the 128-bit security levels compare quite well [^1] Following [1] we estimated 8 NAND gates for a flip-flop, whereas Trivium estimated 12 NAND gates for a flip-flop. Using 12 NAND gates for a flip-flop, our proposal at the 80-bit security level requires about 2395.5 NAND gates, whereas Trivium requires about 3488 NAND gates. with famous ciphers such as Trivium [9] and Grain-128a [1] which offer 80-bit and 128-bit security respectively. For the other security levels, we are not aware of other stream ciphers which have such low gate counts. So our revival of the nonlinear filter model of stream ciphers leads to concrete proposals which offer a combination of both provable security against well known classes of attacks at a desired level of security and also low gate count.
@@
We note that there are some old (prior to the advent of algebraic attacks) works [50, 51, 30] on efficient implementation of Boolean functions on a large number of variables targeted towards the nonlinear combiner model of stream ciphers. There are also a few later works [45, 18] on implementation of Maiorana-McFarland type functions. These works, however, do not cover the implementation of the functions that we consider, the reason being that these functions themselves do not appear earlier in the literature.
@@ 4 Concrete Proposals
We revisit the filter generator model where the state of an LFSR of length $L$ is filtered using a Boolean function. By $\mathcal{S}(L,m)$, we denote the class of stream ciphers obtained from an LFSR of length $L$ with a primitive connection polynomial, and with $f_{2m+1}$ as the filtering function. For the $\kappa$-bit security level, we assume that the stream cipher supports a $\kappa$-bit secret key and a $\kappa$-bit initialisation vector (IV), and so $L\geq 2\kappa$. Suppose that the state $(s_{L-1},\ldots,s_{0})$ of the LFSR leads to the next state $(s_{L},\ldots,s_{1})$, i.e. the state is right shifted once and the next bit $s_{L}$ becomes the leftmost bit. The initial state of the LFSR is constructed as $(k_{\kappa-1},\ldots,k_{0},v_{\kappa-1},\ldots,v_{0},b_{1},\ldots,b_{L-2\kappa})$, where $(k_{\kappa-1},\ldots,k_{0})$ is the key, $(v_{\kappa-1},\ldots,v_{0})$ is the IV, and $(b_{1},\ldots,b_{L-2\kappa})\neq\mathbf{0}_{L-2\kappa}$ is an arbitrary constant bit string.
We next specify the tap positions, i.e. which state bits of the LFSR are fed into the filtering function. From the state $(s_{L-1},\ldots,s_{0})$ of the LFSR, the value of $W$ in $f_{2m+1}(W,\mathbf{X},\mathbf{Y})$ is the state bit $s_{L-\kappa+m}$, the values of the variables in $\mathbf{X}$ are the state bits $s_{L-\kappa+m-1},\ldots,s_{L-\kappa}$, and the values of the variables in $\mathbf{Y}$ are the state bits $s_{L-2\kappa+m-1},\ldots,s_{L-2\kappa}$. So if we write the state vector $\mathbf{s}=(s_{L-1},\ldots,s_{0})$ as $\mathbf{s}_{1}$, $\mathbf{s}_{2}$ and $\mathbf{s}_{3}$, where $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ are $\kappa$-bit strings and $\mathbf{s}_{3}$ is a $(L-2\kappa)$-bit string, then the tap positions for $W$ and $\mathbf{X}$ are the rightmost $m+1$ bits of $\mathbf{s}_{1}$, and the tap positions for $\mathbf{Y}$ are the rightmost tap positions for $\mathbf{s}_{2}$. Note that the tap positions are divided into two clusters, the first cluster consisting of $m+1$ consecutive tap positions, the second cluster consisting of $m$ tap positions, and there is a gap of $\kappa-m$ positions between the two clusters. Our choice of tap positions does not follow the recommendation given in the literature [27, 28, 33, 46, 58]. Nonetheless, we later argue that the design can resist state guessing attacks. This is due to the fact that we use Boolean functions on many more variables than what has been proposed in the literature. Let $M$ be the length of the interval between the first tap position for $\mathbf{X}$ and the last tap position for $\mathbf{Y}$, including the end points. Then $M=\kappa+m$. The quantity $M$ will be later required to quantify resistance to inversion attacks [27].
We assume that the stream cipher goes through an initialisation phase of $2\kappa$ steps, where no keystream bit is produced, and instead the keystream bit is fed back by XORing it with the next bit of the LFSR. So after the initialisation phase the leftmost $2\kappa$ bits of the LFSR are complex nonlinear functions of the key and the IV. Keystream generation starts after the initialisation phase.
@@
The parameters $L$ and $m$ are the design parameters of the stream cipher. The goal is to choose these parameters so as to resist well known classes of known plaintext attacks which result in recovering the key. We consider two basic attack parameters, namely the number $N$ of keystream bits that is required for a successful attack, and the time complexity $T$ of the attack. To ensure security at level $\kappa$ any attack should require $T>2^{\kappa}$. The condition $L\geq 2\kappa$ ensures that the size of the state is at least twice the size of the key. This prevents certain (theoretical) time/memory trade-off attacks [34]. We assume that at most $2^{B}$ keystream bits are generated from a single key and IV pair. So $N$ is at most $2^{B}$, as otherwise the attack cannot be mounted. Further, a basic condition is that $N\leq T$, i.e. some operation is performed on each keystream bit that is required. So for $\kappa$-bit security using $2^{B}$ keystream bits, we have $B\leq\kappa$. In the analysis below, our goal is to determine values of $L$ and $m$ such that the stream cipher is not vulnerable to the known types of attacks. For this concrete analysis, we ignore small constants appearing in the expressions for $N$ and $T$ (i.e. we consider these constants to be 1) and focus only on the exponential components of these quantities.
@@
Anderson leakage.
An interesting method for exploiting leakage by the filtering function was introduced in [3]. In this approach, the focus is on knowing how much information is leaked by the filtering function even if the input LFSR sequence is replaced by a purely random bit sequence. Examples where such leakage can be observed were provided in [3] for the case of filtering functions on a few variables. We call such leakage to be Anderson leakage. Avoiding Anderson leakage amounts to showing that if the input sequence (to the filtering function) is purely random then the keystream sequence is also purely random. Theorem 2 in [27] provides a sufficient condition to ensure this property. The theorem states that if the $n$-variable filtering function is of the form $Z_{1}+g(Z_{2},\ldots,Z_{n})$, then there is no Anderson leakage. Since the filtering function $f$ (built from ${\sf MM}_{2m+1}$) that we propose is of the stated form, it does not exhibit Anderson leakage.
State guessing attacks.
These are guess-then-determine attacks. The idea is to guess some bits and then use the obtained keystream to verify the guess. The first attack of this type was described in [27] and is called the inversion attack. The inversion attack is specifically applicable to filter functions of the form $X\oplus g(\mathbf{Z})$. Since the filter function that we use is indeed of this form, we need to consider the resistance of the stream cipher to the inversion attack. Consider the bit sequence generated by the LFSR to be $s_{0},s_{1},s_{2},\ldots$ and the corresponding keystream bits be $k_{0},k_{1},k_{2},\ldots$. Due to the placement of the tap positions, the $i$-th keystream bit $k_{i}$ is determined by the bits $s_{i},s_{i-1},\ldots,s_{i-M}$ (see above for the definition of $M$). Further, we have $k_{i}=s_{i}\oplus g(s_{i-1},\ldots,s_{i-M})$, where $g$ is defined from ${\sf MM}_{2m}$ as a function of $M$ variables which is degenerate on the $\kappa-m$ variables. So we have $s_{i}=k_{i}\oplus g(s_{i-1},\ldots,s_{i-M})$, i.e. the $i$-th state bit is determined by the known keystream bit $k_{i}$ and the values of the bits $s_{i-1},\ldots,s_{i-M}$. Once $s_{i}$ is obtained, $s_{i+1}$ is obtained as $s_{i+1}=k_{i+1}\oplus g(s_{i},\ldots,s_{i-M+1})$, similarly, from $s_{i+1}$, one obtains $s_{i+2}$ and so on. In other words, if for some $i$, the values of $s_{i-1},\ldots,s_{i-M}$ are known, then the LFSR sequence can be computed in the forward direction. The attack proceeds by guessing the values of $s_{i-1},\ldots,s_{i-M}$, computing the LFSR sequence for $L$ steps to determine the state of the LFSR and then using this state to generate a keystream, if the generated keystream matches with the obtained keystream, then the guess is correct. Since there are $M$ bits whose values have to be guessed, the method requires $2^{M}$ guesses and about $Lm$ operations for each guess. For ${\sf S}(L,m)$, the value of $M$ is $\kappa+m$. So the inversion attack requires more than $2^{\kappa+m}$ operations on ${\sf S}(L,m)$. This proves the resistance of ${\sf S}(L,m)$ to the inversion attack. Note that the inversion attack is prevented due to $M$ being sufficiently large. This was one of the countermeasures to inversion attack that was already proposed in [27].
The inversion attack was later extended to the generalised inversion attack, the filter state guessing attack, and the generalised filter state guessing attack [27, 28, 33, 46, 58]. As one of the countermeasures to these attacks, some design criteria, such as full positive difference set, for choosing tap positions have been suggested. Choosing the tap positions to be consecutive has been mentioned to be the worst possible choice [33]. Our choice of tap positions consists of two clusters where the tap positions in each cluster are consecutive. So based on the recommendations in the literature, our choice of tap positions is close to the worst possible design choice. Nonetheless, we argue that $\mathcal{S}(L,m)$ resists all kinds of state guessing attacks. The reason is that we use a filtering function on many more variables than has been considered in the literature. All the recommendations in the literature implicitly assume that the number of variables of the filtering function is much smaller than the target security level $\kappa$. Since we use filtering functions with number of variables close to $\kappa$, the recommendations on tap positions in the literature do not apply to our case.
Let $n$ be the number of variables of the filtering function. At any point of time, each of the input bits to the filtering function can be written as a linear function of the $L$ initial state bits of the LFSR. So knowing the values of the $n$ input bits at any point of time provide $n$ linear equations in $L$ variables. If the values of the input bits at $c$ points of time are known, where $nc\geq L$, then one obtains a system of $L$ linear equations in $L$ variables and hence can solve this system to obtain the initial state of the LFSR. Of course, one does not know the input bits to the filtering function at any point of time. There are $2^{n}$ possible values of these input bits. The keystream bits $k_{1},k_{2},\ldots$, i.e. the outputs of the filtering function are known. Since the filtering function is balanced, observing one bit of the keystream, say $k_{i}$, rules out half of the possible values of the corresponding input bits to the filtering function, i.e. there are $2^{n-1}$ possible values of the input bits that give rise to the observed keystream bit. So guessing one of these $2^{n-1}$ values provides $n$ linear equations in $L$ variables. An input bit to the filtering function at time point $i$ may also be an input bit to the filtering function at a few other time points. For $\mathcal{S}(L,m)$ the two sets of input bits to the filtering function at two successive time points $i$ and $i+1$ have $12m-$ 1 bits in common. Suppose we have guessed the values of the input bits for time point $i$. Then the number of possible values for the input bits at time point $i+1$ is $4$. Observing $k_{i+1}$ and again using the fact that the filtering function is balanced, the number of possible values for the input bits at time point $i+1$ reduces to $2$. So the observed keystream bits $k_{i}$ and $k_{i+1}$ together restrict the number of possibilites of values of the input bits at time point $i$ and at time point $i+1$ to be $2^{n}$. By guessing one of these $2^{n}$ values, we obtain a system of $2n$ linear equations in $L$ variables. Extending this idea, we see that by guessing one of $2^{n+c-1}$ values of the $c$ consecutive values of the input bits at time points $i,i+1,\ldots,i+c-1$, we obtain a system of $nc$ equations in $L$ variables. If $n+c>L$, then we can solve the linear system and obtain the initial state of the LFSR. Solving the linear system requires $L^{3}$ operations. Since a linear system has to be solved for each guess, the complexity of the attack is $2^{n+c-1}L^{3}$. Suppose the following condition holds.
$2^{n+c-1}L^{3}$ $>$ $2^{\kappa}\mbox{ where $c$ is the least positive integer such that $n+c>L$.}$ (9)
Then no form of the state guessing attack succeeds against $\mathcal{S}(L,m)$ at the $\kappa$-bit security level.
In the above analysis, we saw that the tap positions of $\mathcal{S}(L,m)$ (two clusters, with consecutive positions in each cluster) result in the fact that once the values of the input bits to the $i$-th position is guessed, there are only two possible values for the input bits to the $(i+1)$-st position. This is due to the fact of the huge overlap between the sets of input bits in the $i$-th and $(i+1)$-st positions. The recommendations for tap positions in the literature [27, 28, 33, 46, 58] aim to reduce this overlap. A commonly used recommendation is full positive difference set, i.e. the absolute values of the differences between the tap positions should be distinct. For a filtering function with $n$ variables, this recommendation results in the gap between the first and the last tap positions to be more than $21+2+\cdots+(n-1)=n(n-1)/$ 2 (since the successive differences must be distinct). So the size of the LFSR is at least quadratic in $n$. From efficiency considerations, this forces the value of $n$ to be small. By not following this recommendation, we have done away with the condition that $L$ must be at least quadratic in $n$. Instead, by considering the fundamental requirement behind state guessing attacks, we identified (9) as the condition to resist such attacks. This allows use to choose $n$ to be quite large and close to $\kappa$. Of course, this is possible due to the fact that we have an efficient method for implementing the filtering functions on such large values of $n$.